# XIFBench: Evaluating Large Language Models on Multilingual Instruction Following

### v1

@@ A.4 Multilingual Expansion Details
Your goal is to ensure all essential requirements from the original instruction are preserved in the translation, allowing for minor linguistic variations.
@@
paper Translated Instruction:
   - The machine-translated version of the "original instruction" in {translated_language}.
   - It is the instruction needing to be evaluated for quality.
@@
   - The English machine-translated version of the "translated instruction".
   - It serves as an additional tool to verify the translation’s accuracy.
@@
   - Definition: YES/NO questions designed to assess whether a response follows the instruction. Each question corresponds to an explicit and atomic requirement from the original instruction.
   - Purpose: During evaluation, each question is used to locate the corresponding portions in the instruction that need to be evaluated in the translation.
   - Metadata: Information such as "content requirement - include content" indicates the category and direction of the requirement, aiding in identifying the portion to be evaluated.
   - Implicit Requirements: Occasionally, there may be implicit requirements not explicitly stated in the instruction. These should be identified and treated as exceptions (detailed below) to ensure they do not affect the overall validation.
@@ 1. Portion Identification:
   - Extract the minimal relevant portion from "original instruction", "translated instruction", and "back-translated instruction" using the "evaluation requirements" as a guide.
   - If the instruction combines multiple requirements, isolate only the portion relevant to the current requirement to ensure focused evaluation.
   - For implicit requirements, replace all those 3 portions with the placeholder "[IMPLICIT_REQUIREMENT]".
@@ 2. Observation: Provide a concise observation focusing on:
   - How well the "translated portion" in "translated instruction" preserves the semantics of the "original portion" in "original instruction"?
   - Use the "back-translated portion" in "back-translated instruction" to verify the accuracy of the translation.
   - For implicit requirements, explain why the requirement is not applicable.
@@
   - For implicit requirements, always assign a score of "0/5".
@@
{Provide a concise observation on semantic preservation and accuracy. For implicit requirements, explain why the requirement is implicit.}
@@
### Original Instruction
{original_instruction}
### Translated Instruction (in {translated_language})
{translated_instruction}
### Back-translated Instruction
{back_translated_instruction}
### Evaluation Requirements
{evaluation_requirements}
### Your Validations
@@
In this section, we detail the translation validation process introduced in Section 3.1.4. Since our annotators are not native speakers of Russian, Arabic, Hindi, or Swahili, we leverage Google Translate and GPT-4o to automatically assess translation quality. Our validation procedure follows a requirement-based evaluation framework, ensuring that each instruction and its translation are analyzed for both consistency and accuracy.
Specifically, for each instruction-translation pair, we prompt GPT-4o to identify corresponding segments in the English instruction and its translated counterpart, using predefined evaluation requirements as reference points. Additionally, we incorporate back-translation via Google Translate to mitigate potential biases, particularly for low-resource languages. The LLM annotator then provides a concise assessment of translation correctness and assigns a quality score on a 1-5 scale, where a score of 3 serves as the threshold for acceptable translation quality. The detailed prompt used for this evaluation is provided in Table 14.
@@
5 93.46 95.09 93.26 94.31 89.53
4 5.30 3.29 4.38 4.30 7.60
3 0.93 1.01 0.97 0.81 1.63
2 0.27 0.54 0.85 0.39 0.97
1 0.04 0.08 0.54 0.19 0.27
Table 15: Distribution of translation quality scores across different languages (%).
To quantify translation quality, we analyze the distribution of evaluation requirement scores across different languages. The results, summarized in Table 15, show that the majority of translations receive high-quality ratings. Notably, the proportion of requirements scoring below 3 remains low across all languages, with Arabic exhibiting the highest at just 1.39%. These findings suggest that the translations are generally accurate and consistent with the original English instructions, and their inclusion in the benchmark does not significantly impact the overall evaluation results.
Appendix B Dataset Statistics
As illustrated in Figure 6, the distribution of evaluation requirements in the Easy Set is largely dominated by Content constraints, reflecting a lack of diversity in constraint types. However, despite being constructed based on the Easy Set, the Hard Set exhibits a more balanced distribution across all categories, indicating a diverse and well-balanced constraint augmentation process.
Figure 6: Requirement category distribution across the Easy Set and Hard Set.
Further analysis of the requirement count distribution across the Easy and Hard Sets is shown in Figure 7. While both sets follow a similar uniform distribution, the Hard Set exhibits a broader range of requirement counts, with a pronounced peak at 5 requirements, compared to 2 in the Easy Set. This distribution highlights the greater diversity and complexity of the Hard Set, making it a more challenging evaluation benchmark for LLMs.
Figure 7: Requirement count distribution across the Easy Set and Hard Set.
Appendix C Evaluation Protocol Details
@@
   - The English instruction given to a Large Language Model (LLM).
   - It also serves as the source for deriving evaluation requirements.
@@
   - The LLM’s response to the English instruction.
   - It will be evaluated against each requirement.
@@
   - The YES/NO questions designed to assess whether the response follows the instruction. Each question corresponds to an explicit and atomic requirement from the English instruction.
   - Additional metadata (e.g., "content requirement - include content") provides context about the type and focus of the requirement. Use this metadata to identify the specific aspect of the response that needs to be evaluated.
@@
   - The {translated_language} instruction given to a Large Language Model (LLM).
   - The directive part is translated from the English instruction, while other parts remain unchanged intentionally.
@@
   - The LLM’s response to the {translated_language} instruction.
   - It will be evaluated against each requirement.
@@
   - The original English instruction from which the {translated_language} instruction is translated. Unchanged parts are omitted for brevity.
   - It serves as the source for deriving evaluation requirements and helps in understanding the {translated_language} instruction.
@@
   - The YES/NO questions designed to assess whether the response follows the instruction. Each question corresponds to an explicit and atomic requirement from the English instruction.
   - Additional metadata (e.g., "content requirement - include content") provides context about the type and focus of the requirement. Use this metadata to identify the specific aspect of the response that needs to be evaluated.
@@
Appendix D Experiment Details
D.1 Automatic Evaluation
The official API names of the models used in Section 4.1 are: gpt-4o-2024-11-20, gemini-2.0-flash-001 and claude-3-5-sonnet-20241022. The open-source models and their repositories on Hugging Face [^6] https://huggingface.co/models are: THUDM/glm-4-9b-chat, meta-llama/Llama-3.1-8B-Instruct, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-7B-Instruct, Qwen/Qwen2.5-14B-Instruct, and Qwen/Qwen2.5-72B-Instruct. Evaluating an LLM on the entire XIFBench dataset using gpt-4o-2024-08-06 (without batching) costs approximately \$24.
D.2 Agreement Evaluation
@@
   - note: (Optional) Provide a brief reason if you disagree with the evaluator’s decision. Important Notes:
   - The initial true or false values in do_agree are placeholders – they should not influence your judgment.
   - While the note field is optional, it is strongly recommended when you disagree, as it helps us to analyze disagreement cases.
@@
   - false: Select false if the response fails to satisfy the requirement, contains inaccuracies, omits relevant details, or introduces ambiguity. This strict evaluation means that even if the response is mostly correct, partially correct, or correct under certain conditions, it should still be evaluated as false.
@@
   - style requirement - specify writing style: Is the response written in an academic style?
@@
   - situation requirement - assume role: Does the response adopt the perspective of a historian?
@@
   - situation requirement - establish purpose: Does the response aim to persuade the reader?
@@
   - format requirement - specify output format/pattern: Is the response formatted as a checklist?
@@
   - numerical requirement - limit length parameters: Does the response have exactly five sentences?
@@ 2. For each requirement in instruction_following_evaluation, follow these steps:
   - Read the requirement and observation to understand the context.
   - Read the [PROMPT] and [RESPONSE] sections in the TXT file to examine the actual prompt and response.
   - Use Google Translate or other tools if needed to understand the [PROMPT] and [RESPONSE] content.
   - Determine if the [RESPONSE] fully meets the requirement.
   - The observation can help navigate the response, but do not rely on it as evidence for agreement or disagreement, as it may not always accurately reflect the response.

### v2

@@ C.4 Multilingual Expansion Details
Your goal is to ensure all essential requirements from the original instruction are preserved in the translation.
@@
Translated Instruction:
- The machine-translated version of the "original instruction" in {translated_language}.
- It is the instruction needing to be evaluated for quality.
@@
- The English machine-translated version of the "translated instruction".
- It serves as an additional tool to verify the translation’s accuracy.
@@
- Definition: YES/NO questions designed to assess whether a response follows the instruction. Each question corresponds to an explicit and atomic requirement from the original instruction.
- Purpose: During evaluation, each question is used to locate the corresponding portions in the translation that need to be evaluated.
- Metadata: Information such as "content requirement - include content" indicates the category and direction of the requirement, aiding in identifying the portion to be evaluated.
- Implicit Requirements: Occasionally, there may be implicit requirements not explicitly stated in the instruction. These should be identified and treated as exceptions (detailed below) to ensure they do not affect the overall validation.
@@ 1. Portion Identification:
- Extract the minimal relevant portion from "original instruction", "translated instruction", and "back-translated instruction" using the "evaluation requirements" as a guide.
- If the instruction combines multiple requirements, isolate only the portion relevant to the current requirement to ensure focusing.
- For implicit requirements, replace all those 3 portions with the placeholder "[IMPLICIT_REQUIREMENT]".
@@ 2. Observation: Provide a concise observation focusing on:
- How well the "translated portion" in "translated instruction" preserves the semantics of the "original portion" in "original instruction"?
- Use the "back-translated portion" in "back-translated instruction" to verify the accuracy of the translation.
- For implicit requirements, explain why the requirement is not applicable.
@@
- For implicit requirements, always assign a score of "0/5".
@@
{Provide a concise observation on semantic preservation and accuracy. For implicit requirements, explain why.}
@@
### Original Instruction ### Back-translated Instruction {original_instruction} {back_translated_instruction} ### Translated Instruction (in {translated_language}) ### Evaluation Requirements {translated_instruction} {evaluation_requirements} ### Your Validations
@@
In this section, we detail the translation validation process introduced in Section 3.1.4. Since manual annotation of the entire dataset for Chinese, Russian, Arabic, Hindi, and Swahili is costly and time-consuming, we leverage GPT-4o and Google Translate to automatically assess translation quality. Our validation procedure follows a requirement-based evaluation framework, ensuring that each instruction and its translation are analyzed for both accuracy and consistency.
Specifically, for each instruction-translation pair, we prompt GPT-4o to identify corresponding segments in the English instruction and its translated counterpart, using predefined evaluation requirements as reference points. Additionally, we incorporate back-translation via Google Translate to mitigate potential biases, particularly for low-resource languages. The LLM annotator then provides a concise assessment of translation accuracy and assigns a quality score on a 1-5 scale (1 = Failed Preservation, 5 = Perfect Preservation), where a score of 3 serves as the threshold for acceptable translation quality. The evaluation prompt is provided in Table 14.
@@
5 93.5 95.1 93.3 94.3 89.5
4 5.3 3.3 4.4 4.3 7.6
3 0.9 1.0 1.0 0.8 1.6
2 0.3 0.5 0.9 0.4 1.0
1 0.0 0.1 0.5 0.2 0.3
(a) Distribution of translation quality scores (%) from GPT-4o on the full dataset.
Score Zh Ru Ar Hi Sw
5 81.5 79.5 83.0 76.8 79.3
4 13.3 15.9 12.0 17.8 15.9
3 4.2 3.3 2.7 4.3 3.7
2 1.0 0.8 2.3 0.8 1.0
1 0.0 0.4 0.0 0.2 0.0
(b) Distribution of translation quality scores (%) from human annotation on 20% of the dataset.
Figure 6: Comparison of translation quality score distributions across languages. Left: GPT-4o annotations on the full dataset. Right: Human annotations on a 10% subset.
To quantify translation quality, we analyze the distribution of evaluation requirement scores across different languages. To further validate the automatic evaluation and assess its agreement with human judgments, we randomly sample 20% of English instructions (18 from the Easy Set and 90 from the Hard Set) and their translations in all non-English languages. For each language, two native-speaking annotators independently scores the sampled translations on a 1-5 scale, following guidelines slightly adapted from Table 14 to facilitate human annotation. Additionally, the segments extracted by GPT-4o during scoring are provided to annotators to help identify the relevant portions for each requirement. We take the lower of the two scores for each translation to conservatively estimate translation quality.
The results, summarized in Table 6(a), show that the vast majority of translations receive high-quality ratings. Notably, the proportion of requirements scoring below 3 is low across all languages, with Arabic exhibiting the highest at only 1.4%. These findings suggest that the translations are largely accurate and consistent with the original English instructions, making them suitable for inclusion in the benchmark without significantly affecting overall evaluation outcomes.
Human annotation results on the sampled subset, shown in Table 6(b), further support this conclusion: over 97.7% of all translations are rated $\geq$3. Disagreements occur in about 25% of segments, but are mostly minor (e.g., human = 4 vs. GPT-4o = 5), indicating that GPT-4o may slightly overestimate quality. These results validate our automatic scoring as a reasonable proxy for translation quality and support the reliability of our multilingual benchmark.
Appendix D Dataset Statistics Details
As illustrated in Figure 7(a), the distribution of evaluation requirements in the Easy Set is largely dominated by Content constraints, reflecting a lack of diversity in constraint types. However, despite being constructed based on the Easy Set, the Hard Set exhibits a more balanced distribution across all categories, indicating a diverse and well-balanced constraint augmentation process.
Further analysis of the requirement count distribution across the Easy and Hard Sets is shown in Figure 7(b). While both sets follow a similar uniform distribution, the Hard Set exhibits a broader range of requirement counts, with a pronounced peak at 5 requirements, compared to 2 in the Easy Set. This distribution highlights the greater diversity and complexity of the Hard Set, making it a more challenging evaluation benchmark for LLMs.
(a) Requirement category distribution across the Easy Set and Hard Set.
(b) Requirement count distribution across the Easy Set and Hard Set.
Figure 7: Distribution analysis of requirements in the dataset. Left: Category distribution across Easy and Hard sets. Right: Count distribution across Easy and Hard sets.
Appendix E Evaluation Protocol Details
@@
- The English instruction given to a Large Language Model (LLM).
- It also serves as the source for deriving evaluation requirements.
@@
- The LLM’s response to the English instruction.
- It will be evaluated against each requirement.
@@
- The YES/NO questions designed to assess whether the response follows the instruction. Each question corresponds to an explicit and atomic requirement from the English instruction.
- Additional metadata (e.g., "content requirement - include content") provides context about the type and focus of the requirement. Use this metadata to identify the specific aspect of the response that needs to be evaluated.
@@
- The {translated_language} instruction given to a Large Language Model (LLM).
- The directive part is translated from the English instruction, while other parts remain unchanged intentionally.
@@
- The LLM’s response to the {translated_language} instruction.
- It will be evaluated against each requirement.
@@
- The original English instruction from which the {translated_language} instruction is translated. Unchanged parts are omitted for brevity.
- It serves as the source for deriving evaluation requirements and helps in understanding the {translated_language} instruction.
@@
- The YES/NO questions designed to assess whether the response follows the instruction. Each question corresponds to an explicit and atomic requirement from the English instruction.
- Additional metadata (e.g., "content requirement - include content") provides context about the type and focus of the requirement. Use this metadata to identify the specific aspect of the response that needs to be evaluated.
@@
Appendix F Experiment Details
In this section, we provide detailed information about the experiments discussed in Section 4, including supplemental details on the automatic evaluation of LLMs, and detailed description of agreement between GPT-4o and human annotations.
F.1 Automatic Evaluation
The official API names of the models used in Section 4.1 are: gpt-4o-2024-11-20, gemini-2.0-flash-001 and claude-3-5-sonnet-20241022. The open-source models and their repositories on Hugging Face [^5] https://huggingface.co/models are: THUDM/glm-4-9b-chat, meta-llama/Llama-3.1-8B-Instruct, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-7B-Instruct, Qwen/Qwen2.5-14B-Instruct, and Qwen/Qwen2.5-72B-Instruct. For the inference of open-source models, we used at most 2 H20-80G GPUs for inference. Evaluating an LLM on the entire XIFBench dataset using gpt-4o-2024-08-06 (without batching) costs approximately \$24.
F.2 Agreement Evaluation
@@
- note: (Optional) Provide a brief reason if you disagree with the evaluator’s decision. Important Notes:
- The initial true or false values in do_agree are placeholders – they should not influence your judgment.
- While the note field is optional, it is strongly recommended when you disagree, as it helps us to analyze disagreement cases.
@@
- false: Select false if the response fails to satisfy the requirement, contains inaccuracies, omits relevant details, or introduces ambiguity. This strict evaluation means that even if the response is mostly correct, partially correct, or correct under certain conditions, it should still be evaluated as false.
@@
- style requirement - specify writing style: Is the response written in an academic style?
@@
- situation requirement - assume role: Does the response adopt the perspective of a historian?
@@
- situation requirement - establish purpose: Does the response aim to persuade the reader?
@@
- format requirement - specify output format/pattern: Is the response formatted as a checklist?
@@
- numerical requirement - limit length parameters: Does the response have exactly five sentences?
@@ 2. For each requirement in instruction_following_evaluation, follow these steps:
- Read the requirement and observation to understand the context.
- Read the [PROMPT] and [RESPONSE] sections in the TXT file to examine the actual prompt and response.
- Use Google Translate or other tools if needed to understand the [PROMPT] and [RESPONSE] content.
- Determine if the [RESPONSE] fully meets the requirement.
- The observation can help navigate the response, but do not rely on it as evidence for agreement or disagreement, as it may not always accurately reflect the response.
@@
In this section, we detail the process of evaluating agreement between GPT-4o’s assessments and human annotations across the three evaluation settings outlined in Section 4.2.
  • Direct Scoring (DS): GPT-4o scores each response on a scale from 1 to 10. This approach does not explicitly assess whether each constraint is satisfied, but rather offers a holistic judgment of the response quality. Following prior work [11, 12], responses scoring above 5 are considered to meet all evaluation requirements for calculating agreement against our human-labeled requirement-level ground truth. We adopt the evaluation prompt from Zheng et al. [10] without modification.
  • Ours w/ Trans. Reqs. (Ours with Translated Requirements): GPT-4o evaluates responses using evaluation requirements translated into the same language as the instruction. For example, a Chinese instruction-response pair was evaluated with Chinese-translated evaluation requirements and metadata (categories, dimensions). This setting simulates a multilingual evaluation scenario without a shared semantic anchor, an issue raised in Section 3.3, allowing us to examine how the language of evaluation criteria affects reliability and consistency.
  • Ours w/ Eng. Reqs. (Ours with English Requirements): This is our proposed evaluation protocol. GPT-4o evaluates all instruction-response pairs using the original English evaluation requirements, regardless of the language of the instruction. For instance, a Swahili instruction-response pair is evaluated using English evaluation requirements and metadata. This ensures a consistent semantic anchor across languages, avoiding translation-induced inaccuracies and ensuring cross-linguistic comparability.
We use Google Translate to obtain translated requirements and translated metadata for the Ours w/ Trans. Reqs. setting. The evaluation prompt template (see Table 15) is adapted by replacing language-specific parts—such as the descriptive text "English instruction" and its corresponding placeholder {english_instruction}—with their language-specific equivalents (e.g., "Chinese instruction" and {chinese_instruction}).